# Co-VisiON: Co-Visibility ReasONing on Sparse Image Sets of Indoor Scenes

### v1

@@ G.2 Zero-shot Evaluation
G.3 Qualitative Result of Mask Prediction
@@
H Failure mode analysis
We add more baselines: Gemini-2.0-Flash [57] (a recent VLM with better spatial understanding), SpatialRGPT [11] (a recent Spatial-VLM), Covis-freeze (a variant with the CrocoV2 backbone frozen except for the last two layers) and SigLIP (a contrastive learning baseline). All baselines were evaluated on 100 Gibson test scenarios (5 per scene), except for Gemini-2.0-Flash, which was run on 20 scenarios using the first scenario from each scene due to limited time and compute during rebuttal. A full-scale benchmark for Gemini-2.0-Flash will be made available upon paper acceptance. Although we cannot and were not meant to be exhaustive, we added the above mentioned more recent baselines in our camera-ready version because of their better spatial reasoning capabilities. For the GPT-4o and Gemini 2.0 Flash baselines, we provide image pairs along with a prompt that explicitly asks the model to reason about co-visibility between the two images. SpatialRGPT, is built on Vila[31], which supports multi-image input. We feed image pairs into its visual encoder and use the fine-tuned LLaMA3-70B to reason about co-visibility from the encoded tokens. We use the same prompt for GPT-4o, Gemini 2.0 Flash, and SpatialRGPT, as shown inFig. V. For the failure mode analysis, we categorize test scenes into easy, medium, and hard levels based on scene sparsity and pairwise overlap (see the caption of Tab. III).
Figure V: An example of the prompt used in our experiments. A unified prompt is adopted across all vision-language models, including GPT-4o, Gemini-2.0-Flash, and SpatialRGPT.
Table III: Best IoU performance across different difficulty levels under two evaluation protocols: (1) Image Overlap, categorized based on the overlap ratio of each image pair — easy: $\geq 50\%$, medium: $\geq 10\%$ and $<50\%$, hard: $<10\%$. (2) Scene Sparsity, defined by the average pairwise overlap within each scene — easy: $\geq 10\%$, medium: $\geq 4\%$ and $<10\%$, hard: $<4\%$. Top-3 performers in each column are highlighted using the color scheme consistent with Table 2 in the main paper.
Setting Method Easy Med. Hard Avg.
Image Overlap (edge level) GPT-4o 0.97 0.92 0.34 0.63
Gemini-2.0-Flash 1.00 0.99 0.14 0.42
Qwen2.5-VL 72b 1.00 0.99 0.14 0.41
SigLIP 1.00 1.00 0.13 0.35
SpatialRGPT 1.00 0.99 0.13 0.35
Covis 0.99 0.88 0.24 0.59
Covis (freeze) 1.00 0.89 0.30 0.61
Scene Sparsity (graph level) GPT-4o 0.83 0.64 0.57 0.63
Gemini-2.0-Flash 0.75 0.39 0.28 0.42
Qwen2.5-VL 72b 0.77 0.40 0.28 0.41
SigLIP 0.73 0.38 0.26 0.35
SpatialRGPT 0.72 0.38 0.26 0.35
Covis 0.80 0.63 0.52 0.59
Covis (freeze) 0.81 0.65 0.54 0.61
Discussion. In Tab. III, GPT-4o outperforms other VLMs. Thanks to R3’s suggestion, Covis (freeze) slightly improves over the original. Both Covis and GPT-4o achieve the most significant improvements over all baselines in hard case scenarios. These hard cases highlight that limited viewpoint overlap and scene sparsity are the primary sources of difficulty.
I Sim2Real
In this experiment, our aim is to demonstrate that the Co-VisiON dataset exhibits a small domain gap compared to real-world environments, such as the AVD dataset [3]. We have tested Covis, ViT, VGG, Resnet, Contrastive, and NetVlad methods on the AVD dataset using pretrained model from Co-VisiON dataset. We observe AUC metric results in Tab. IV to be comparable to those in Tab. 2. Similarly, from Fig. VI, we can see that the co-visibility graph predicted using the Covis closely resembles the manually labelled topology graph.
@@ I Sim2Real
Figure VI: Comparison of manually labeled and Covis predicted Topologies. For clarity, we show 10 sampled images to provide an illustrative example.
Table IV: Zero-shot comparison of AUC values (%) for baseline models pretrained on the Co-VisiON dataset and evaluated on real-world data.

### v3

@@ G.1 Zero-shot Evaluation
G.2 Qualitative Result of Mask Prediction
@@
H Sim2Real
In this experiment, our aim is to demonstrate that the Co-VisiON dataset exhibits a small domain gap compared to real-world environments, such as the AVD dataset [4]. We have tested Covis, ViT, VGG, Resnet, Contrastive, and NetVlad methods on the AVD dataset using pretrained model from Co-VisiON dataset. We observe AUC metric results in Tab. II to be comparable to those in Tab. 4. Similarly, from Fig. V, we can see that the co-visibility graph predicted using the Covis closely resembles the manually labelled topology graph.
@@ H Sim2Real
Figure V: Comparison of manually labeled and Covis predicted Topologies. For clarity, we show 10 sampled images to provide an illustrative example.
Table II: Zero-shot comparison of AUC values (%) for baseline models pretrained on the Co-VisiON dataset and evaluated on real-world data.